OpenStudy (anonymous):

The perimeter of a retangle is 40m. The length is 8m more than twice the width. find the demensions. what is the lenght and width please?

OpenStudy (anonymous):

it is 16 lenght and 4 width?

OpenStudy (anonymous):

Let p= perimeter, l=length, and w=width. p=2l+2w=40 l=8+2w 40=2(8+2w) +2w Divide both sides by two: 20=8+2w+w=8+3w Subtract 8: 12=3w w=12/3=4

OpenStudy (anonymous):

don't forget your units!

OpenStudy (anonymous):

So what does that make the width and what does that make the length? please I will give a medal!!!

OpenStudy (anonymous):

p = 2(w + h) 40 = 2(4 + h) 40 = 8 + 2h 32 = 2h h = 16

OpenStudy (anonymous):

I only need to the width and the length!

OpenStudy (anonymous):

W = 4, L = 16

OpenStudy (anonymous):

that is what I put first!